# Ronny Seiger5, Agnes Koschmider2, Stefanie Rinderle-Ma4, Barbara Weber5 and Estefania Serral 7

### v2

@@ 3.4.3 CAIRO
The CAIRO metamodel, shown in Figure 5, has been proposed in (franceschetti2023event), motivated by the need for an IoT-driven process monitoring and conformance checking across a wide range of processes and domains. The CAIRO metamodel aims at being agnostic with respect to process and domain characteristics such as level of process structuredness, process modelling paradigm (e.g., imperative or declarative), level of automation of the process activities, and process-awareness, i.e., the support from a Process-aware information system for orchestrating process executions. The foundational assumptions are the observability of process events in terms of IoT events, the knowledge of the monitored process (either formally or informally specified), and a defined mapping of low-level IoT events into high-level process events. At the core of the metamodel is the notion of CAIRO event, i.e., an event that is contextualized, Atomic, Instantaneous, Relevant, and Observable. The contextualized property is represented in the metamodel by the association between the CAIRO event class and the Event attribute class, which posits that event attributes associate contextual information to an event. The property of being atomic (i.e., events have all-or-nothing semantics) is enforced in the metamodel by cardinality constraints associating each instance of an Event occurrence of a CAIRO event with exactly one instance of a Process event from a known process specification. The property of being instantaneous binds a CAIRO event to a specific point in time, and is realized by associating a CAIRO event with exactly one timestamp attribute. The property of being relevant requires an event to refer to a relevant state change of a process, and is enforced in the metamodel by associating the CAIRO event class with the Monitoring point class, which models applications detecting activity occurrences based on IoT events. The property of being observable is realized in the CAIRO metamodel with the association between class Monitoring point and class Expression, representing the aggregation of the readings from instances of the associated class Sensor.
Figure 5: CAIRO metamodel (franceschetti2023event)
@@ 4 Methodology
The proposed CORE metamodel is the artifact resulting from the application of the design science research methodology described in (Hevner.2004_DSRinIS). According to this methodology, the development of the artifact represented as the CORE metamodel follows a sequence of activities. These activities are based on the Design Science Research Methodology (DSRM) process model by (Peffers.2008_DSRM_ProcessModel), which consists of six activities: 1) Identify Problem and Motivate, 2) Define Objectives of a Solution, 3) Design and Development of an Artifact, 4) Demonstration, 5) Evaluation, and 6) Communication. This manuscript serves as the medium of Communication of the developed artifact. The remaining activities are described in the following sections, along with insights on how they have been conducted in the context of the development of the CORE metamodel .
@@ 4.1 Identify Problem and Motivate
The development of a common metamodel is motivated by the need to integrate IoT data with process event data in IoT-enhanced business processes–a task complicated by the heterogeneity, noise, and low granularity of IoT sensor data (bertrand2022assessing). As shown in Sect. 3, current models differ significantly in their assumptions, application domains, and design objectives, resulting in fragmentation that impedes interoperability and data reuse. A common metamodel aims to bridge these gaps by capturing core concepts shared across existing models, facilitating translation and standardization of logs across diverse use cases. This promotes broader adoption by fostering the exchange of IoT-enhanced process data.
@@ 4.2 Define Objectives of a Solution
    – Different Granularity Levels: Different granularity levels must be representable, including a) IoT events (e.g., low-level sensor data, higher level IoT events (seiger2023data)) and b) process events (e.g., state changes in workflow executions).
@@
    – Traceability: Traceability across different granularity levels is essential. Therefore, it should be possible a) to establish relationships between low-level and high-level events, such as linking IoT events to process events. Furthermore, since event abstraction can be part of functions or algorithms, b) it should be possible to annotate an event with the abstraction algorithm or expression used, indicating that it was generated through this process (GonzalezLopezdeMurillas2019).
@@
    – Semantic Annotations: In semantics-aware process analytics (rebmann2022enabling), prior knowledge can be used, if available. Therefore, the metamodel should be able to represent a) procedural knowledge, such as information about the process model (formal or informal, e.g., in natural language), b) the relationship between sensors and the process within the infrastructure, and c) prior knowledge from other data sources, such as databases or taxonomies.
@@
    – Flexible Case Notion: For some use-cases, a single case notion is not sufficient. Therefore, a more flexible case notion is desired, which allows to not only link events to one specific case but to (potentially multiple) entities being involved in its execution. This enables analysis from different points of view (i.e., based on different corresponding entities such as machines, locations, etc. (GonzalezLopezdeMurillas2019).
@@
    – Different Data Sources: Data in an IoT environment can originate from different sources, which are very heterogeneous. It should be possible to integrate and represent various data sources, including a) information systems (IS), such as ERP systems or databases containing process metadata, b) sensors, which provide real-time data like temperature, humidity, or motion, and c) actuators, which record actions such as opening a valve, starting a motor, or adjusting a device’s settings (GonzalezLopezdeMurillas2019). Moreover, it should be possible to represent these sources in a unified manner, despite the high level of heterogeneity.
@@ 4.3 Design and Development of an Artifact
The development of the CORE metamodel required a structured and scientifically grounded alignment process among international experts from six specialized working groups. To achieve this, a participatory and iterative development approach has been adopted, drawing on the principles of Participatory Design (bodker2022participatory) and iterative prototyping (camburn2017design). This methodology ensured the active participation of all stakeholders while allowing continuous validation and refinement of the model.
@@ 5.2 Metamodel Presentation
  • General objects: other types of objects, which do not fall into one of the two previously introduced classes, are explicitly mentioned for emphasis. Examples include subprocesses or activity objects (benzin2024inexa).
@@ 5.2.1 Event Type Definitions
In process mining, an event is commonly defined as the execution of (part of) an activity gunther2014xes. However, an event has a very different meaning in the IoT literature, where it is often seen as an observation made by a sensor janowicz2019sosa; Bertrand2024Feb. To resolve this discrepancy, our model introduces two types of events, IoT events and process events, and gives them explicit definitions.
An IoT event represents a change in a physical phenomenon, which is typically derived from one observation of one IoT device such as a sensor, e.g., a measurement made by a temperature sensor. To account for complex physical phenomena, which cannot directly be tracked by an IoT device, but require to process the observations made by one IoT device or to aggregate data collected by multiple IoT devices, we also make it possible to derive a higher-level IoT event from one or more lower-level IoT events seiger2023data, which can be characterized by a label indicating the semantics of the event. An example of such an IoT event is the detection of a peak in time series sensor data, which requires processing consecutive observations of one sensor using an algorithm.
@@ 5.2.2 Objects Taxonomy
First, the Case objects represent traditional objects as defined by the OCED (fahland2024towards) and OCEL standards (berti2024ocel). Such objects are entities that are involved in the execution of activities. As such, they represent cases in subprocesses that provide different views on long end-to-end processes. Examples include orders, product bottles, batches, etc.
@@ 6.1 Architectural Overview
The decision of an architecture for CORE metamodel adheres to specific requirements (e.g., model conformity, rapid prototyping). Model conformity ensures correctness and consistency. Model-driven engineering (MDE) allows quick iterations due to frequent metamodel changes. Given the dynamic CORE metamodel development, we prioritize simplicity, reliability, and testability as key requirements for the architecture decision. Based on these requirements and following established best practices (ford2020fundamentals), we decided to use a Layered Software Architecture as an architectural style for our implementation. Following this, we define three layers: the Input Layer is necessary to parse the different existing logs (see Section˜3.4). The Data Transformation Layer links concepts between the input model and CORE metamodel (see Section˜5.1 and Figure˜2). Finally, the OCEL Representation Layer transforms and represents our model in OCEL (see Figure˜6). The architecture is visualized in Figure˜10, consisting of the three layers, Input Layer, Data Transformation Layer, OCEL Representation Layer. To closely reflect our implementation, we explicitly include Python-specific type annotations in the UML representations.
@@ 6.4 OCEL Representation Layer
The implementation should be independent of a specific OCEL implementation (e.g., pm4py.OCEL [(pm4py)], ocpa [(ocpa)]) and follow the structure provided in Figure˜2. This abstraction not only enhances the maintainability and flexibility of the implementation, but also ensures that the output remains reusable and interoperable.
@@ 7.1.1 Use case description
In this use case, we describe how we transcribe a log originally stored in the NICE format to our common CORE metamodel . This synthetic event log, first presented in (bertrand2023nice), represents the behavior of two inhabitants living in a smart home for four weeks. To generate it, we used the smart spaces simulator presented in (veneruso2023model), which we adapted to generate NICE logs. More specifically, the two users were programmed to execute daily life activities such as Cook and eat, Do the dishes, Drink, Eat cold, Eat warm, Exercise, Go work, Sleep, Rest, Shop, Use Computer, Watch tv and Wc. Moreover, the simulator was programmed so that during one week, the users show a different behavior, i.e., they stay at home the whole day instead of going to work in the morning and coming back home in the evening during weekdays. This change in behavior is due to extremely high outside temperatures during that week, which are tracked by temperature sensors and logged as IoT events.
@@
  • Flexibility of the process: human routines are characterized by their context dependency, their lack of structure, the repetition of activities and the execution of multiple activities in parallel (Leotta2015ApplyingPM);
@@ 7.2.1 Use Case Description
In manufacturing processes, multiple (IoT) devices operate collaboratively to produce a part. These processes encompass various steps, including the handling of parts, which may involve multiple machines, transportation between different machines, and the measuring of parts. In the following, two use cases are presented that demonstrate the integration of IoT devices into manufacturing processes. In each case, the processes are managed by a process execution engine and are logged using the XES DataStream extension, ensuring real-time tracking and analysis of production events. This approach facilitates enhanced process monitoring, optimization, and traceability, contributing to improved efficiency and product quality. (datastream):
  • In the IoT lab at Trier University [^5] More information can be found at https://iot.uni-trier.de., a Fischertechnik learning factory is used to carry out research related to BPM (Malburg.2020_BPMAndIoT). The use case includes machines which are equipped with several types of sensors that collect production-related data, such as the state of connected light barriers or switches. Complementing this Internet of Things (IoT) data, information pertaining to process executions is systematically recorded and maintained as event logs, facilitating comprehensive monitoring and analysis of manufacturing processes (Grueger.2023).
@@ 7.2.2 Translation to core log format
The logs for the use cases described before are logged using the XES DataStream extension (datastream). In the following, it is described how they can be translated into the core model format:
@@ 7.3.1 Use case description
In this use case, the focus is on the IoT-supported online monitoring and conformance checking of the blood donation process, augmented with hand hygiene guidelines from the World Health Organization (WHO) that prescribe additional process steps of hand disinfection (franceschetti2023proambition). The goal is to be able to provide a timely feedback to the healthcare workers regarding the conformance with the hand hygiene requirements, reducing the risk of contamination that could threaten the patients’ health. As the process is fully manual and executed without the support of a process engine, the challenge here is to be able to generate in real-time IoT events from the process execution and abstract them into the respective process events. However, ambiguities deriving from multiple process events such as Apply tourniquet and Insert needle being represented by the same IoT events pose a challenge to the event abstraction. This results in multiple process events being generated from a given set of IoT events, with attached a quantification of the respective ambiguity.
@@ 8 Discussion
  • Flexibility and Extensibility: By supporting semantic annotations, metadata representation, and diverse data sources, the model accommodates a wide range of use cases that require the representation of additional data beyond events, as demonstrated in the evaluation in Sect. 7. This extensible architecture paves the way for future extensions and customized adaptations, for example capturing domain-specific knowledge. We expect this to be instrumental to enable further advancements in context-aware as well as semantics-aware process mining (rebmann2022enabling).
  • Interoperability: We deliberately made the choice of implementing the CORE metamodel within the OCEL 2.0 format (berti2024ocel). Currently, OCEL 2.0 is the go-to format adopted by researchers in process mining (goossens2024object). Therefore, we expect that this will facilitate uptake by researchers and integration with existing use cases, fostering data exchange and reuse.
@@
Overall, the CORE metamodel addresses several challenges, providing benefits for researchers and practitioners. Still, we recognize some limitations of the metamodel. Specifically, the reconciliation of diverse requirements from different use cases and research groups resulted in a number of trade-offs in the expressive power of the metamodel. This results in the lack of explicit concepts in the CORE metamodel able to represent use case-specific concepts, such as the Formal Model concept from CAIRO, which represents a formal process specification and is relevant in the smart healthcare use case. Nevertheless, the CORE metamodel remains open for specialization and extension to ensure that also use-case specific requirements can be met. This can be achieved, for example, by leveraging event attributes or by specializing the general object concept and its associations (cf. Sect. 5). This extensibility can be supported by a model-driven engineering (MDE) (kent2002model) approach, where the CORE metamodel acts as a central specification from which new artifacts can be derived. With this approach, adopting the metamodel and introducing new subclasses or domain-specific concepts does not require a full re-implementation, as changes can be integrated (semi-)automatically using MDE tools. We emphasize that these trade-offs reflect deliberate design choices made during the metamodel development to prioritize generalizability and interoperability–which we consider major goals of the metamodel.

### v3

@@ 3.4.3 CAIRO
The CAIRO metamodel, shown in Figure 5, has been proposed in (Franceschetti et al. 2023c), motivated by the need for an IoT-driven process monitoring and conformance checking across a wide range of processes and domains. The CAIRO metamodel aims at being agnostic with respect to process and domain characteristics such as level of process structuredness, process modelling paradigm (e.g., imperative or declarative), level of automation of the process activities, and process-awareness, i.e., the support from a Process-aware information system for orchestrating process executions. The foundational assumptions are the observability of process events in terms of IoT events, the knowledge of the monitored process (either formally or informally specified), and a defined mapping of low-level IoT events into high-level process events. At the core of the metamodel is the notion of CAIRO event, i.e., an event that is contextualized, Atomic, Instantaneous, Relevant, and Observable. The contextualized property is represented in the metamodel by the association between the CAIRO event class and the Event attribute class, which posits that event attributes associate contextual information to an event. The property of being atomic (i.e., events have all-or-nothing semantics) is enforced in the metamodel by cardinality constraints associating each instance of an Event occurrence of a CAIRO event with exactly one instance of a Process event from a known process specification. The property of being instantaneous binds a CAIRO event to a specific point in time, and is realized by associating a CAIRO event with exactly one timestamp attribute. The property of being relevant requires an event to refer to a relevant state change of a process, and is enforced in the metamodel by associating the CAIRO event class with the Monitoring point class, which models applications detecting activity occurrences based on IoT events. The property of being observable is realized in the CAIRO metamodel with the association between class Monitoring point and class Expression, representing the aggregation of the readings from instances of the associated class Sensor.
Figure 5: CAIRO metamodel (Franceschetti et al. 2023c)
@@ 4 Methodology
The proposed CORE metamodel is the artifact resulting from the application of the design science research methodology described in (Hevner et al. 2004). According to this methodology, the development of the artifact represented as the CORE metamodel follows a sequence of activities. These activities are based on the Design Science Research Methodology (DSRM) process model by (Peffers et al. 2007), which consists of six activities: 1) Identify Problem and Motivate, 2) Define Objectives of a Solution, 3) Design and Development of an Artifact, 4) Demonstration, 5) Evaluation, and 6) Communication. This manuscript serves as the medium of Communication of the developed artifact. The remaining activities are described in the following sections, along with insights on how they have been conducted in the context of the development of the CORE metamodel .
@@ 4.1 Identify Problem and Motivate
The development of a common metamodel is motivated by the need to integrate IoT data with process event data in IoT-enhanced business processes–a task complicated by the heterogeneity, noise, and low granularity of IoT sensor data (Bertrand et al. 2022). As shown in Sect. 3, current models differ significantly in their assumptions, application domains, and design objectives, resulting in fragmentation that impedes interoperability and data reuse. A common metamodel aims to bridge these gaps by capturing core concepts shared across existing models, facilitating translation and standardization of logs across diverse use cases. This promotes broader adoption by fostering the exchange of IoT-enhanced process data.
@@ 4.2 Define Objectives of a Solution
    – Different Granularity Levels: Different granularity levels must be representable, including a) IoT events (e.g., low-level sensor data, higher level IoT events (Seiger et al. 2023)) and b) process events (e.g., state changes in workflow executions).
@@
    – Traceability: Traceability across different granularity levels is essential. Therefore, it should be possible a) to establish relationships between low-level and high-level events, such as linking IoT events to process events. Furthermore, since event abstraction can be part of functions or algorithms, b) it should be possible to annotate an event with the abstraction algorithm or expression used, indicating that it was generated through this process (de Murillas et al. 2019).
@@
    – Semantic Annotations: In semantics-aware process analytics (Rebmann and van der Aa 2022), prior knowledge can be used, if available. Therefore, the metamodel should be able to represent a) procedural knowledge, such as information about the process model (formal or informal, e.g., in natural language), b) the relationship between sensors and the process within the infrastructure, and c) prior knowledge from other data sources, such as databases or taxonomies.
@@
    – Flexible Case Notion: For some use-cases, a single case notion is not sufficient. Therefore, a more flexible case notion is desired, which allows to not only link events to one specific case but to (potentially multiple) entities being involved in its execution. This enables analysis from different points of view (i.e., based on different corresponding entities such as machines, locations, etc. (de Murillas et al. 2019).
@@
    – Different Data Sources: Data in an IoT environment can originate from different sources, which are very heterogeneous. It should be possible to integrate and represent various data sources, including a) information systems (IS), such as ERP systems or databases containing process metadata, b) sensors, which provide real-time data like temperature, humidity, or motion, and c) actuators, which record actions such as opening a valve, starting a motor, or adjusting a device’s settings (de Murillas et al. 2019). Moreover, it should be possible to represent these sources in a unified manner, despite the high level of heterogeneity.
@@ 4.3 Design and Development of an Artifact
The development of the CORE metamodel required a structured and scientifically grounded alignment process among international experts from six specialized working groups. To achieve this, a participatory and iterative development approach has been adopted, drawing on the principles of Participatory Design (Bødker et al. 2022) and iterative prototyping (Camburn et al. 2017). This methodology ensured the active participation of all stakeholders while allowing continuous validation and refinement of the model.
@@ 5.2 Metamodel Presentation
  • General objects: other types of objects, which do not fall into one of the two previously introduced classes, are explicitly mentioned for emphasis. Examples include subprocesses or activity objects (Benzin et al. 2024).
@@ 5.2.1 Event Type Definitions
In process mining, an event is commonly defined as the execution of (part of) an activity Günther and Verbeek (2014). However, an event has a very different meaning in the IoT literature, where it is often seen as an observation made by a sensor Janowicz et al. (2019); Bertrand et al. (2024). To resolve this discrepancy, our model introduces two types of events, IoT events and process events, and gives them explicit definitions.
An IoT event represents a change in a physical phenomenon, which is typically derived from one observation of one IoT device such as a sensor, e.g., a measurement made by a temperature sensor. To account for complex physical phenomena, which cannot directly be tracked by an IoT device, but require to process the observations made by one IoT device or to aggregate data collected by multiple IoT devices, we also make it possible to derive a higher-level IoT event from one or more lower-level IoT events Seiger et al. (2023), which can be characterized by a label indicating the semantics of the event. An example of such an IoT event is the detection of a peak in time series sensor data, which requires processing consecutive observations of one sensor using an algorithm.
@@ 5.2.2 Objects Taxonomy
First, the Case objects represent traditional objects as defined by the OCED (Fahland et al. 2024) and OCEL standards (Berti et al. 2024). Such objects are entities that are involved in the execution of activities. As such, they represent cases in subprocesses that provide different views on long end-to-end processes. Examples include orders, product bottles, batches, etc.
@@ 6.1 Architectural Overview
The decision of an architecture for CORE metamodel adheres to specific requirements (e.g., model conformity, rapid prototyping). Model conformity ensures correctness and consistency. Model-driven engineering (MDE) allows quick iterations due to frequent metamodel changes. Given the dynamic CORE metamodel development, we prioritize simplicity, reliability, and testability as key requirements for the architecture decision. Based on these requirements and following established best practices (Richards and Ford 2020), we decided to use a Layered Software Architecture as an architectural style for our implementation. Following this, we define three layers: the Input Layer is necessary to parse the different existing logs (see Section˜3.4). The Data Transformation Layer links concepts between the input model and CORE metamodel (see Section˜5.1 and Figure˜2). Finally, the OCEL Representation Layer transforms and represents our model in OCEL (see Figure˜6). The architecture is visualized in Figure˜10, consisting of the three layers, Input Layer, Data Transformation Layer, OCEL Representation Layer. To closely reflect our implementation, we explicitly include Python-specific type annotations in the UML representations.
@@ 6.4 OCEL Representation Layer
The implementation should be independent of a specific OCEL implementation (e.g., pm4py.OCEL [(Berti et al. 2023)], ocpa [(Adams et al. 2022)]) and follow the structure provided in Figure˜2. This abstraction not only enhances the maintainability and flexibility of the implementation, but also ensures that the output remains reusable and interoperable.
@@ 7.1.1 Use case description
In this use case, we describe how we transcribe a log originally stored in the NICE format to our common CORE metamodel . This synthetic event log, first presented in (Bertrand et al. 2026), represents the behavior of two inhabitants living in a smart home for four weeks. To generate it, we used the smart spaces simulator presented in (Veneruso et al. 2023), which we adapted to generate NICE logs. More specifically, the two users were programmed to execute daily life activities such as Cook and eat, Do the dishes, Drink, Eat cold, Eat warm, Exercise, Go work, Sleep, Rest, Shop, Use Computer, Watch tv and Wc. Moreover, the simulator was programmed so that during one week, the users show a different behavior, i.e., they stay at home the whole day instead of going to work in the morning and coming back home in the evening during weekdays. This change in behavior is due to extremely high outside temperatures during that week, which are tracked by temperature sensors and logged as IoT events.
@@
  • Flexibility of the process: human routines are characterized by their context dependency, their lack of structure, the repetition of activities and the execution of multiple activities in parallel (Leotta et al. 2015);
@@ 7.2.1 Use Case Description
In manufacturing processes, multiple (IoT) devices operate collaboratively to produce a part. These processes encompass various steps, including the handling of parts, which may involve multiple machines, transportation between different machines, and the measuring of parts. In the following, two use cases are presented that demonstrate the integration of IoT devices into manufacturing processes. In each case, the processes are managed by a process execution engine and are logged using the XES DataStream extension, ensuring real-time tracking and analysis of production events. This approach facilitates enhanced process monitoring, optimization, and traceability, contributing to improved efficiency and product quality. (Mangler et al. 2023):
  • In the IoT lab at Trier University [^5] More information can be found at https://iot.uni-trier.de., a Fischertechnik learning factory is used to carry out research related to BPM (Malburg et al. 2020). The use case includes machines which are equipped with several types of sensors that collect production-related data, such as the state of connected light barriers or switches. Complementing this Internet of Things (IoT) data, information pertaining to process executions is systematically recorded and maintained as event logs, facilitating comprehensive monitoring and analysis of manufacturing processes (Grüger et al. 2023).
@@ 7.2.2 Translation to core log format
The logs for the use cases described before are logged using the XES DataStream extension (Mangler et al. 2023). In the following, it is described how they can be translated into the core model format:
@@ 7.3.1 Use case description
In this use case, the focus is on the IoT-supported online monitoring and conformance checking of the blood donation process, augmented with hand hygiene guidelines from the World Health Organization (WHO) that prescribe additional process steps of hand disinfection (Franceschetti et al. 2023a). The goal is to be able to provide a timely feedback to the healthcare workers regarding the conformance with the hand hygiene requirements, reducing the risk of contamination that could threaten the patients’ health. As the process is fully manual and executed without the support of a process engine, the challenge here is to be able to generate in real-time IoT events from the process execution and abstract them into the respective process events. However, ambiguities deriving from multiple process events such as Apply tourniquet and Insert needle being represented by the same IoT events pose a challenge to the event abstraction. This results in multiple process events being generated from a given set of IoT events, with attached a quantification of the respective ambiguity.
@@ 8 Discussion
  • Flexibility and Extensibility: By supporting semantic annotations, metadata representation, and diverse data sources, the model accommodates a wide range of use cases that require the representation of additional data beyond events, as demonstrated in the evaluation in Sect. 7. This extensible architecture paves the way for future extensions and customized adaptations, for example capturing domain-specific knowledge. We expect this to be instrumental to enable further advancements in context-aware as well as semantics-aware process mining (Rebmann and van der Aa 2022).
  • Interoperability: We deliberately made the choice of implementing the CORE metamodel within the OCEL 2.0 format (Berti et al. 2024). Currently, OCEL 2.0 is the go-to format adopted by researchers in process mining (Goossens et al. 2024). Therefore, we expect that this will facilitate uptake by researchers and integration with existing use cases, fostering data exchange and reuse.
@@
Overall, the CORE metamodel addresses several challenges, providing benefits for researchers and practitioners. Still, we recognize some limitations of the metamodel. Specifically, the reconciliation of diverse requirements from different use cases and research groups resulted in a number of trade-offs in the expressive power of the metamodel. This results in the lack of explicit concepts in the CORE metamodel able to represent use case-specific concepts, such as the Formal Model concept from CAIRO, which represents a formal process specification and is relevant in the smart healthcare use case. Nevertheless, the CORE metamodel remains open for specialization and extension to ensure that also use-case specific requirements can be met. This can be achieved, for example, by leveraging event attributes or by specializing the general object concept and its associations (cf. Sect. 5). This extensibility can be supported by a model-driven engineering (MDE) (Kent 2002) approach, where the CORE metamodel acts as a central specification from which new artifacts can be derived. With this approach, adopting the metamodel and introducing new subclasses or domain-specific concepts does not require a full re-implementation, as changes can be integrated (semi-)automatically using MDE tools. We emphasize that these trade-offs reflect deliberate design choices made during the metamodel development to prioritize generalizability and interoperability–which we consider major goals of the metamodel.